OpenStudy (anonymous):

you invested of $4000 in two stocks that paid 5% and 7% annual interest. at the end of the year, the totally interest from these investments was $230 . how much was invested at each rate?

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (anonymous):

@LukeBlueFive

OpenStudy (lukebluefive):

The following two equations can be derived from this word problem: 0.05 * x + 0.07 * y = $230 x + y = $4000

OpenStudy (lukebluefive):

This looks like a substitution problem, so let's separate the variables: y = $4000 - x Then substitute this value for y in the first equation, then solve for x: 0.05 * x + 0.07 * ($4000 - x) = $230

OpenStudy (lukebluefive):

This equals: 0.05 * x + $280 - 0.07 * x = $230 Simplifying: $280 - 0.02 * x = $230 Subtracting $280 from each side: -0.02 * x = -$50 Writing -0.02 as a fraction: \[-\frac{2}{100} * x = -$50\] Simplifying: \[-\frac{1}{50} * x = -$50\] Multiply both sides by -50. That will gives you the amount invested in 5% stock. Subtract it from $4000 to get the amount invested in 7% stock.